OpenStudy (anonymous):

how do you simplify 2^-4? I entered it as in on my calculator and got 0.0625 which is wrong to what my program is looking for

OpenStudy (kinggeorge):

\[2^{-4}={1 \over 2^4} =0.0625\]To change a negative exponent to a positive exponent, take the reciprocal.

OpenStudy (anonymous):

ok and I got the 0.0625 but they say its wrong

OpenStudy (anonymous):

maybe "they" are looking for \(\frac{1}{16}\)

OpenStudy (qqstory):

0.0625 is correct?

OpenStudy (anonymous):

yep they wanted the 1/16 thank you I didnt even think about that

OpenStudy (qqstory):

Seit, are you funny or what? 1/16 = 0.0625. god..